nolegirl24:

What are the zeroes of f(x) = x2 + 5x + 6? I need help solving it out

dude:

You can factor or use the quadratic equation, which one feels easier for you?

dude:

Or do you just want me to walk you through one of them, whichever seems easier

nolegirl24:

Can you walk me through it please..

dude:

Factoring seems easier imo Do you understand the general idea of factoring?

nolegirl24:

Not really

nolegirl24:

I was never able to get how it worked

dude:

Okay, so you want to find 2 numbers that multiply to 6 first What are the multiples of 6?

nolegirl24:

1,6 2,3 3,2

dude:

Right, which of those numbers can add up to 5?

nolegirl24:

2,3 3,2

dude:

2,3 and 3,2 are the same thing you dont have to re-write them Basically your equation is just \((x+first~multiple)(x+second~multiple)\) = \((x+2)(x+3)\)

dude:

Does that make sense?

nolegirl24:

Yeah, I was just writing it down cause I will end up forgetting easily and I am studying for my final exam.

dude:

Ah okay, also that isnt your final answer, when it asks for `zeroes` always note that you have to set them equal to 0 \(x+2=0\) and \(x+3=0\) Solve for x on each \(x+2\color{red}{-2}=0\color{red}{-2}\) and \(x+3\color{red}{-3}=0\color{red}{-3}\) x=-2 and x=-3 are your zeroes

nolegirl24:

So my answer would be x = -2,-3?

dude:

Yes

nolegirl24:

Thank you, I understand it better now ^.^

dude:

Ofcs ofcs